Volume 2, Number 4 October 1993

TEX and TUG NEWS

for and by the TEX community

A Publication of the TEX Users Group Electronic version TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font selection scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@tug.org with the subject line NEWSLETTER. Advertisements are accepted in TTN: contact the TUG office for details.

TEX Users Group  Phone: +1 (805) 963-1338
P.O. Box 869 FAX: +1 (805) 963-8358
Santa Barbara, CA 93102 USA E-mail: tug@tug.org

TEX is a trademark of the American Mathematical Society.

TEX and TUG NEWS (ISSN 1065-240X) is published quarterly by the TEX Users Group, 735 State Street, P.O. Box 869, Santa Barbara, CA 93102, U.S.A. The 1993 dues for individual members are $60.00 for regular members and $30.00 for students.

Second-class postage paid at Santa Barbara, CA, and additional mailing offices. Postmaster: Send address changes to TEX and TUG NEWS, TEX Users Group, P.O. Box 869, Santa Barbara, CA 93102, U.S.A.

Membership in the TEX Users Group is for the calendar year, and includes all issues of TUGboat and TEX and TUG NEWS for the year in which membership begins or is renewed. Individual membership is open only to named individuals, and carries with it such rights and responsibilities as voting in the annual election.

Editorial

Another year is coming to a close, and plans for 1994 have to begin. We have a new board coming on stream in January (see p. 20); sites for future TUG meetings are being sought (p. 23); and we hope to see you amongst the list of new or renewed members in the TEX Users Group (1994 rates remain at ’93 levels). Just before we leave 1993 … 31 December 1993 is the new deadline for the A-in-L^{A}TEX contest (p. 14), so do send in your entries!

But you don’t have to wait till 1994 to use our new e-mail address! In September, tug@tug.org officially came on-line. The office can be reached at:

TUG Office

                    E-mail:  tug@tug.org

Phone number: 805-963-1338
Fax number:  805-963-8358

A correction for an item which I’d noted in my editorial of the last issue: TEXhelp: The On-Line TEX Handbook is a joint effort by both Arvind Borde and Tom Rokicki. I only listed Arvind as author — my apologies for the oversight, Tom. Along with the correction, Arvind sent the following additional information (see also p. 3):

… it’s a hypertext guide to TEX, L^{A}TEX and AMS-TEX that allows you to look up descriptions and definitions of commands on-line; it also contains discussions of typesetting topics. The initial release will run only on DOS machines, but MS-Windows and Unix versions are to follow. We’re hoping that this electronic guide will be found a useful tool by people who want to write TEX but who don’t have a printed manual on hand, or who want to look things up *fast*.

Something very interesting happened earlier this month. The Goudy Center, Rochester Institute of Technology, has been in touch with TUG about courses which may be of interest to our members (see Upcoming Events, p. 25). And it seems to me that there is potential for a two-way interchange of expertise between this center, which is in the School of Printing Management and Sciences, and our members, with their knowledge and experience with TEX. We’ll keep you posted on future developments.

I’m putting together some information on using TEX for linguistics typesetting. I’d like to get in touch with others who have some expertise with fonts, writing specialised macros, encoding files with all those wonderful “design features” which pepper linguistics papers: tree diagrams, glosses, the phonetic alphabet (and any others), pseudo-mathematical presentation of rules and principles, and so on. There will be a meeting of the Linguistic Society of America in Boston the first week of January, 1994. It is possible that the conference organisers might be interested in a display of how TEX can be used in the field of linguistics. As well, we may be able to produce a useful little guide on the subject. If you’d like to join forces (and maybe find a solution or two yourself!), drop me a line at: cthiele@ccs.carleton.ca, or call me at +1 613-727-0427 (Canada).

The UK-TUG is planning a few meetings over the next few months. As we went to press, no firm dates or locations had been set, but if you’re in the UK, there will be a meeting in either November or December, and there’s also a 2-day L^{A}TEX meeting planned for the spring. Stay tuned both to this newsletter, and the Calendar section in upcoming issues of TUGboat.

Recently, Don Hosek polled readers of the TYPO-L list about a new magazine on typography he’d like to launch. The magazine will now definitely go ahead. For more information, write to Don at 555 Guilford, Claremont, California 91711, or FAX him at +1 909-625-1342.

Electronic distribution of TTN will now be taken care of by Michael Doob; our thanks go to Chris Carruthers for having gotten the distribution started.

Christina Thiele

Editor, TEX and TUG NEWS

Positions Available

TEX and TUG NEWS needs a new Editor, as well as a Columnist for the Upcoming Events page.

The Editor should have experience with L^{A}TEX, which is what TTN uses; the editor should also be familiar with electronic submissions and correspondence, in order to coordinate and manage the work done by contributors and the team of columnists already in place. The editor should have a broad awareness of TEX and TUG activities, be able to seek out potential news items, and have a strong sense of duty (!), even though this is an entirely volunteer effort. An extensive network of contacts is a definite asset.

The Upcoming Events Columnist gathers information on activities related to TUG, to other TEX user groups, and items which might interest readers of TTN. The columnist then submits the finished column to the editor.

If you are interested in either position, please contact the current editor

at cthiele@ccs.carleton.ca.

New Publications

Peter Schmitt

                                         schmitt@awirap.bitnet
                           a8131dal@awiuni11.edvz.univie.ac.at

Arvind Borde and Tomas Rokicki. A User’s Guide for TEXHelp; The On-Line TEX Handbook. Academic Press, 1993. v + 28pp. ISBN 0-12-1176040-1.

• This small booklet documents the TEXHelp program for IBM PC and compatible microcomputers.

Arvind Borde. Mathematical TEX by Example.

Academic Press, 1993. 356pp. US$19.95. ISBN 0-12-117645-2.

• As the title suggests, this book is an introduction to TEX by means of examples. They almost entirely consist of pairs of pages: one contains some piece of TEX code (with explanations and comments added via footnotes), while the other shows the output produced by this code.

Nicholas J. Higham. Handbook of Writing for the Mathematical Sciences. SIAM — Society for Industrial and Applied Mathematics, Philadelphia, 1992. xii + 241pp. US$21.50. ISBN 0-89871-314-5.

• This book is intended for all authors of technical papers (and books) or talks, and covers virtually all relevant issues. The author discusses questions of style as well as practical matters. The use of computers in the publishing process — including TEX, of course — is discussed in detail. An appendix contains a list of TEX and L^{A}TEX symbols. Among the various useful materials are hints for Internet use and an extensive bibliography.

Christian Rolland. LATEX guide pratique.

Addison-Wesley France, July 1993. 280pp. approx. 190 FF. ISBN 2-87908-025-0. (Distributed by Bordas, Bordas code 009808.)

Latest issues of other TEX newsletters

GUST: Zeszyt 2 (1993), 42pp. Official newsletter of the Polish User Group, Grupa Użytkowników Systemu TEX.

La Lettre GUTenberg: no. 1, juillet 1993, 16pp.

Die TEXnische Komödie: 5. Jahrgang, Heft 1, Mai 1993, 72pp; Heft 2, Sep. 1993,

TUGboat: vol. 14, no. 1, April 1993, 92pp; vol. 14, no. 2, July 1993, 64pp.

Typographer’s Inn

Peter Flynn

University College Cork

                                               pflynn@curia.ucc.ie

Interletter spacing

Anyone who reads any of the TEX-related mailing lists or Usenet news groups must be well aware that the Computer Modern (cm) fonts are much criticised for alleged ugliness. cm-bashing can be very much in vogue at times, but most of its critics still allow that at 1270 dpi and above, it is extremely easy on the eye to read. This is largely due to the proportions of the interletter spacing, which are wider than in many faces used in other dtp systems.

In most visual dtp systems you can tweak the interletter space (sometimes called the ‘tracking’). This is regarded as inadvisable for plain text, as type designers go to considerable trouble to adjust interletter spacing for optimal readability, yet it looks as if Adobe has narrowed it in most of the fonts they distribute.

At its worst, there are even systems, including some wordprocessors, which actually pride themselves on their ability to achieve unhyphenated justified setting by slackening the interletter spacing. While this is acceptable in newspapers, with their exceptionally narrow columns, it looks very poor in normal text work.

One day I will have to find out if it is possible to bodge the afm2tfm parameters to generate a decent appearance for the Adobe fonts, but I’m not sure how this will come out once it’s passed through dvips and handed to a typesetter. Does anyone have any experience in this?

Interword spacing

Richard Southall gave a provocative talk at Aston about plain TEX’s limitations in the matter of justification. In hand composition, interword spacing is initially a fixed space while the line is being set in the stick. Justification proceeds by the addition or substitution of fixed increments selected from ens, thicks, mids, thins and hairspaces, whereas TEX uses infinitely variable spacing (almost: the increments are tiny). At one stage the trend in the uk was for very close setting, sometimes with no space at all between a fullpoint at the end of a word and the following capital of the next sentence, a practice which caused much debate in my typography classes when I was a student.

It is true that properly done, hand-set matter is more pleasing to read because of the care and attention which is paid by the skilled compositor. TEX, being only a program, cannot hope to achieve this level of sophistication, but the interesting point for the development of the New Typesetting System (nts) is the possibility of being able to switch to a system of justification based on discrete increments. Unless, of course, someone would like to implement this in TEX as it stands…

Poetry setting

There was only one entry for my challenge to find a way of automating the folding of long lines into the space left above or below by a preceding or following short measure. Unfortunately the solution arrived in my mailbox while I was actually at Aston, so pints were not distributed, but the winner was Jonathan Kew, whose complex but elegant macros are available by ftp from curia.ucc.ie in

pub/tex/VerseMacs.tex and VerseTest.tex — well done Jonathan!

Dropped caps

A final note on the macros for dropped initial capitals: I was reading Anne McCaffrey’s Crystal Line in the us paperback edition (Del Rey/Ballantine, 1992, ISBN 0–345–38491–1) and was amused to see the opening of the first chapter with a 3–line dropped cap for the test below. Quite correctly, all three opening quotes and the A were dropped into the paragraph, but this takes up over a third of the measure. Some days you just can’t win…

“ ‘A

nd a star to steer her by,’ ” Killashandra Ree shouted to herself. Not that Lars Dahl could

have heard her over the roar of the sea crashing against the bow of the Angel and the humming tension of the wind through the sail stays and across the sloop’s mainsail.

Mailbox

Yannis Haralambous commented on TEX’s lack of a long ‘s’ for 17th century work, pointing out that he has included it in the new dc fonts, as well as ligaturing it. I can see I shall have to start using the dc fonts now!

Jim Graham asked about whether to break a line before or after an em rule, if one occurs at a breakpoint. I prefer afterwards, always, but Nelson Beebe said his definition permits linebreaks before and after when used in narrow setting (he de-

fines \emrule as \penalty\exhyphenpenalty---\penalty\exhyphenpenalty).

David Carlisle fixed my flow.sty (running text round an illustration) so that it handles indentation correctly, and a new copy is in /pub/tex on

curia.ucc.ie .

Finally a reminder that you can join TYPO-L, the typography mailing list,

by sending sub typo-l forename surname to listserv@irlearn.ucd.ie and

carry on the discussion there.

Thanks to all who contributed — and I didn’t forget to sign off this time.

“Hey — it works!”

Jeremy Gibbons University of Auckland

                                            jeremy@cs.aukuni.ac.nz

Welcome to “Hey — it works!”. This column is designed for little tricks — even ones that seem trivial or inelegant — that others may find useful. If you have any such items, please send them in to me at the address above. I am particularly looking for short items (up to about a page).

We start off with a little note from our noble editor, describing a trick she used in laying out this issue to get one-column material following two-column on the same page. Then there is a piece from David Lau about ‘striking’ (drawing a line) through a piece of text, and a piece from Anita Hoover on making rules of different thicknesses in L^{A}TEX alignments. Finally, I describe a modification to L^{A}TEX’s footnote macro that I used in this issue in order to place verbatim material in footnotes.

One-column material below two-column_{Christina Thiele cthiele@ccs.carleton.ca}

When you get to the end of this issue, we have an index done with the \twocolumn coding, described in TTN 1,3:15–16. My problem was to fit a fullwidth text below the two-column material. I opted for the cheap and sneaky solution: I simply put the text at the end of the second column, and then \hspace’d it leftwards by −6 cm! No argument from TEX, not even a mutter about an overfull \hbox.

   ... end of last index entry in \twocolumn.
\hspace{-6cm}
   \begin{tabular}{||l||}
      first cell of table ... \\
      ... last cell of table
   \end{tabular}
}
\onecolumn

Striking through text _{David Lau Richmond, British Columbia david@worg.questor.org}

I present below a macro, \strikethrough, which will reproduce its argument with a line through it; for example, ‘\underline{strike}’.

\def\strikethrough#1{%
    {%
        \setbox0=\hbox{#1}%
        \dimen0=\ht0 \dimen1=\dp0
        \setbox1=\vbox{
            \box0 \vskip-\dimen1\vskip-0.8ex \hrule
        }%
        \ht1=\dimen0 \dp1=\dimen1
        \box1
    }%
}

This macro works by first typesetting the argument, then backing up vertically by an amount equal to its depth plus 0.8 times the x-height of the current font, and finally overlaying a \hrule on top of the argument.^{1}

Here is the tricky part. It is crucial that the placement of the strikethrough line on the argument not result in a box (\box1) having dimensions different from those of the box (\box0) containing the argument by itself. If the superimposition is done imperfectly, the resultant box will not line up properly with the surrounding text. But according to TEX’s rules of box construction, \box1, as constructed in the code, has a height of \ht0 − 0.8 ex + 0.4 pt and a depth of 0 pt. As a result, the dimensions of \box1 have to be adjusted after it has been constructed, which explains why the line \ht1=\dimen0 \dp1=\dimen1 is necessary.

Note that this macro has a shortcoming: it only works for arguments that are relatively “short”. In the event that lines of text or entire paragraphs are to be “struck through”, a more elaborate approach is called for.

1Another approach is to use horizontal leaders, as in Chapter 21 of the TEXbook:

\def\strikethrough#1{{%

\setbox0=\hbox{#1}% \dimen0 0.8ex\dimen1\dimen0\advance\dimen1 by0.4pt \rlap{\leaders\hrule height \dimen1 depth -\dimen0\hskip\wd0}% \box0 }} This involves backing up horizontally, rather than vertically, and using a rule of negative depth.

Symbols other than the \hrule can be used as leaders. —jg

Thicker \hlines in L^{A}TEX alignments _{Anita Z. Hoover University of Delaware anita@ravel.udel.edu}

In the tabular environment, there is a macro called \hline which produces a horizontal line in a table across all columns. It is defined to be of a certain thickness and can be changed by modifying the dimension specified by \arrayrulewidth. However if you want to keep the original \hline thickness of lines, you need to create a new macro, say \Hline, to produce thicker lines based on a new dimension, say \arrayruleHwidth. You can do this simply by copying the definition for \hline and changing the width of the rules. The definition below shows how; it can be placed in the preamble of your document.^{2}

\newdimen\arrayruleHwidth
\setlength{\arrayruleHwidth}{4pt}
\makeatletter
\def\Hline{\noalign{\ifnum0=`}\fi\hrule \@height \arrayruleHwidth
   \futurelet \@tempa\@xhline}
\makeatother

Alternatively, you could omit the \makeatletter and \makeatother and put the definition in a style file, say Hline.sty.

Here is an example of using the new macro \Hline to produce thicker lines in the tabular environment.

\begin{tabular}{ccc}
   \Hline
   \multicolumn{3}{c}{Very Nice}\\\Hline
   1 & 2 & 3 \\\hline
   4 & 5 & 6 \\\hline
   7 & 8 & 9 \\\Hline
\end{tabular}

This produces

Very Nice 1 2 3 4 5 6 7 8 9

2A similar technique could be used for \vline and \cline. —jg

Footnotes with verbatim material Jeremy Gibbons

There are limitations to what you can do with L^{A}TEX’s \footnote macro; in particular, you cannot include verbatim material (via \verb or \begin{verbatim}) in a footnote. This is because of the way L^{A}TEX’s footnotes are defined; the \footnote macro reads in its entire argument before executing any of it, and the catcodes of the entire argument are fixed when the argument is read — it is not possible to change catcodes within the argument.

On the other hand, plain TEX’s footnote macro (as on page 363 of the TEXbook) does not have this problem; it ‘uses some \bgroup and \futurelet and \aftergroup trickery, so that the footnote text does not need to be a parameter’.

I needed footnotes with verbatim material for Anita’s and David’s items in this issue, in order to put editorial comments about their macros in footnotes. I did this by judiciously coalescing L^{A}TEX’s and plain TEX’s footnote macros, as shown below.

% plainfootnote.sty: incorporate plain TeX's trickery into LaTeX's
% footnote macros, to allow \verb"..." within footnotes
% (the argument to \footnote is not read before it is executed)
\long\def\@footnotetext{%
   \insert\footins\bgroup
      \footnotesize
      \interlinepenalty\interfootnotelinepenalty
      \splittopskip\footnotesep
      \splitmaxdepth \dp\strutbox \floatingpenalty \@MM
      \hsize\columnwidth \@parboxrestore
      \edef\@currentlabel{\csname p@footnote\endcsname\@thefnmark}%
      \@makefntext{\rule{\z@}{\footnotesep}\ignorespaces}%
      \futurelet\next\fo@t
}
\def\fo@t{\ifcat\bgroup\noexpand\next \let\next\f@@t
                                \else \let\next\f@t\fi \next}
\def\f@@t{\bgroup\aftergroup\@foot\let\next}
\def\f@t#1{#1\@foot}
\def\@foot{\strut\egroup}

(L^{A})TEX News

L^{A}TEX2ε — A New Version of L^{A}TEX _{Chris Rowley and Frank Mittelbach L^{A}TEX3 Project}

What is it? L^{A}TEX2ε is a new, standard version of L^{A}TEX — prepared and supported by the L^{A}TEX3 Project Team.

Why is it needed? Over the years many extensions of L^{A}TEX have been developed. This is, of course, a welcome development, since it shows that the L^{A}TEX system is in a healthy state. It has, however, had one unfortunate consequence: there are now several incompatible systems, in the sense of format (.fmt) files, all claiming to be L^{A}TEX. Therefore, in order to process documents coming from various places, a site maintainer needs to provide several format files: L^{A}TEX (with and without NFSS), SliTEX, AMS-L^{A}TEX, and so on. As well, when looking at a source file it is not always clear for which format the document was written.

What will it offer? L^{A}TEX2ε will put an end to this unsatisfactory situation — it will give access to all such extensions based on a single format and thus end the proliferation of mutually incompatible dialects of L^{A}TEX 2.09.

It will use an enhanced version (NFSS2) of the New Font Selection Scheme. Files such as amstex.sty (formerly the AMS-L^{A}TEX format) or slides.sty (formerly the SliTEX format) will become extension packages, all working with this single format.

The introduction of this new version will also make it possible to add a small number of often-requested features (such as extended versions of \newcommand).

To summarize:

• Standardisation: a single format incorporating NFSS2, to replace the present multiplicity of incompatible formats (NFSS, lfonts, pslfonts, etc.)

• Maintenance: a standardised system supported by a reliable maintenance policy.

Compatibilty L^{A}TEX2ε will, as far as possible, adhere to the following principles:

1. Unmodified version 2.09 document files will produce the same output with L^{A}TEX2ε as with version 2.09.

2. All new features of L^{A}TEX2ε will conform to the conventions of version 2.09, making it as easy as possible for current users to learn to use them.

Documentation The new version will be described in a new edition of Leslie Lamport’s L^{A}TEX: A Document Preparation System, and in a new book, The L^{A}TEX Companion by Michael Goossens, Frank Mittelbach and Alexander Samarin (both to be published by Addison-Wesley). The Companion will also contain a complete description of NFSS2.

Distribution Policy Maintenance of the new system will be undertaken by the L^{A}TEX3 Project Team.

A complete distribution of all files, incorporating corrections of errors, will be made available twice a year on fixed dates. This will happen even if there have been no changes to the files, and hence only the release dates have had to be updated. It will therefore be easy to determine whether you have all the latest bug fixes — information which is essential to system maintainers.

Error Reports Error reports can be made using a report generating program latexbug.tex. This will be part of the main distribution.

Error reports will be accepted only if the version of L^{A}TEX2ε that produced the error is not older than one year. Error reports can be sent to this e-mail

address: latex-bugs@rus.uni-stuttgart.de.

Availability The α-test version of L^{A}TEX2ε is scheduled for late fall and the first regular distribution will be in early 1994 (says Frank). This will also contain information and advice to help in the process of updating an installation for use with the new version (says Chris).

L^{A}TEX3 Update: Coming Soon!

There will be a status report on the L^{A}TEX3 Project coming soon. Depending on which publication gets out first, it will either appear in TUGboat 14, #4, or TTN 3,1 (January 1994). Further prototype code for L^{A}TEX3 is taking shape at this very moment (I am informed by reliable sources!). Stay tuned to these pages for more information. It’s coming! –Ch.

Bits and pieces … Christina Thiele

Nelson Beebe sent info on a review of Y&Y’s DVIWindow and DVIPSONE, and comments on outline fonts for TEX: John L. Castri, “Bitmap-free TEX,” Notices of the American Mathematical Society 40, #5 (1993): 464–68.

And from the mailbag … the Y&Y folks have “recently completed the conversion of TEXplorator’s MathTime font set to fully hinted ATM compatible Adobe Type 1 format.” This is version 1.1, “a set of math fonts for use with TEX that goes particularly well with Times Roman for text, but also works well with other text fonts in Adobe Type 1 format.”

You Rang …?

Ed Baker

P.O. Box 642, Norfolk, Mass., 02056 USA

                                                 ejb@world.std.com

[Note: Robert Becker has been sidelined with a severe hand injury. Ed Baker has offered to fill in, and this column is therefore a bit different from the usual fare in “You rang … ?” This time we have a product description. If you have any comments, please send them to Ed at the above address –Ch.]

Virtually all database tools provide the ability to produce plain ASCII reports. Using them with TEX or L^{A}TEX is simply a matter of creating reports that are actual, ready-to-roll, TEX or L^{A}TEX documents.

A popular database product is the Personal Information Manager (PIM). A PIM is essentially a consolidated address book, calendar, phone dialer and mail list manager. They are widely used by people who have to manage large numbers of phone calls, appointments and correspondence.

One of the common commercial PIMs is called ACT (Automated Contact Tracking) produced by Contact Software International. This article will demonstrate how L^{A}TEX can be used directly with ACT to produce a standard business letter.

Each record in ACT is referred to as a contact, containing 74 fields of information. Each field is labeled within the system so that it can be used with the ACT report generator, accessed via a pull-down menu system. The report generator is used to create a master file to be merged with the various contacts. After starting the report generator, simply type in all of the regular L^{A}TEX commands you want to use and insert ACT-replaceable fields by specifying them as shown: “[#FIELD–NAME]”. The attached ACT report is used as a standard business letter shell. It uses the following fields: [#Contact], [#Name], [#Address line 1], [#Address line 2], [#City], [#State], [#ZIP] and [#Salutation].

To use the system, you’d merely select the person you want to write to, use the ACT menu system to select the “Letter” option, and the replaceable fields would be automatically inserted. You could then insert the body of your text, enclosures, cc list and postscript.

The ACT system also provides a function called “Launch” which will start a batch file. You could start L^{A}TEX with this feature, or shell to DOS to run L^{A}TEX and your favorite previewer/printer driver.

The attached report listing illustrates the ACT letter. In this case, all of the

letter parts are defined as \newcommands — \body, \salutation, \postscript,

TEX and TUG NEWS Vol. 2, No. 4, 1993 13

\enclosures, \cclist, and \signatuture. Each \newcommand is placed in the

L^{A}TEX preamble in the inverse order of its likelihood of changing. The contents of the \postscript and \enclosures commands are actually commented out unless you elect to use them.

\documentstyle[art11]{article}
\newcommand{\sendto}{
[#Contact]                      \\
[#Name]                         \\
[#Address_line_1]               \\
[#Address_line_2]               \\
[#City], [#State]  [#ZIP]
}
\newcommand{\salutation}{
Dear [#Salutation],}
\newcommand{\body}{
%%                     -- Your Text Goes Here  --
}
\newcommand{\postscript}{
%% Use the next section for your postscript message
%% remove the two percent signs and continue on the next line
%%{\bf p.s.}\\
}
\newcommand{\enclosures}{\footnotesize
%% Use the next section for your enclosures
%% remove the two percent signs and continue on the next line
%%encl: \\
}
\newcommand{\cclist}{\small
%% remove the two percent signs and put cclist inside parbox
%%CC: \ \parbox[t]{4in}{
%%
%%} % end of parbox
} % end of cclist
\pagestyle{empty}
\setlength{\oddsidemargin}{0pt}
\setlength{\evensidemargin}{0pt}
\setlength{\marginparwidth}{1 in}
\setlength{\marginparsep}{0pt}
\setlength{\topmargin}{0 pt}
\setlength{\headheight}{0pt}
\setlength{\headsep}{0pt}

\setlength{\topskip}{0pt}
\setlength{\footheight}{0pt}
\setlength{\footskip}{0pt}
\setlength{\textheight}{8.5in}
\setlength{\textwidth}{6.0in}
\setlength{\unitlength}{1in}
\newcounter{foo}
\setlength{\parindent}{0pt}
\setlength{\parskip}{2ex}
\newcommand{\signature}{Ed Baker \\
{\em Internet: }{\tt ejb@world.std.com}}
%% End of Preamble %%
\begin{document}
\vspace*{0.75in}
\today \hfill
\sendto\par
\vspace*{5ex}
{\salutation}\par
\vspace*{2ex}
\body\par
\vspace*{2ex}
Sincerely,\par
\vspace*{6ex}
\signature\par
\vspace*{3pt}
\postscript\par
\vspace*{3pt}
\cclist\par
\vfill
\enclosures
\end{document}

A-in-L^{A}TEXContest Deadline Extended

Due to late delivery of many issues of TTN 2,2, which contained contest details, the deadline has been extended to

31 December 1993

See TTN 2,2:28 for contest details.

Reports on Meetings

DANTE’93 in Chemnitz _{Walt A. Obermiller, Stephan Engelke}

DANTE’93, the German TEX meeting, and membership conference of the German TEX users group was held at the Technical University of Chemnitz, March 9–12, 1993. Some 90 registered participants gathered for four days of TEXnical discussions, and to have a good time.

Tutorials The meeting began on Tuesday afternoon, the 9th, with tutorials, a “quaint old tradition” of DANTE meetings. All tutorials were well attended. One of the most popular sessions was with Phil Taylor (“Ask me any question about plain TEX”) which was continued on Friday, after the meeting was supposed to be over. Some tutorials also provided handy booklets, which were very much appreciated.

Membership meeting The DANTE membership meeting was held on Wednesday morning. Friedhelm Sowa (treasurer) gave a financial status report (definitely solvent); then Joachim Lammarsch (president) spoke about DANTE’s activities, and its relationship with other TEX groups. Membership (now at 2,321) has increased steadily. The opening of DANTE’s office in Heidelberg is imminent; the very popular distribution of software and books to members, albeit a lot of work, will move to the new headquarters soon. The quality of maintenance of DANTE’s ftp server in Stuttgart has been improved, and close coordination of activities with the Aston and SHSU servers has been established. Board elections unanimously reconfirmed the old board (Joachim Lammarsch, president; Uwe Untermarzoner, vice president; Friedhelm Sowa, treasurer; and Luzia Dietsche, secretary). The membership meeting was then adjourned, and most of the remaining three days were devoted to the talks on the agenda.

The conference presentations About a dozen papers were presented, covering a wide range of topics, under the general headings of TEX, L^{A}TEX, and , as well as the intersection with PostScript and graphics issues. For L^{A}TEX, there were papers on font problems in a mixed Cyrillic/Roman document (Horst Szillat); macros to create personalized bulk mailings (Thomas Schweikle); a new style file for floats (Hubert Gäßlein); and an account of introducing L^{A}TEX to hospital doctors for writing letters and reports (Bernd-Michael Paschke).

In the TEX area, a description of TEX under Linux (Jens Pönisch); publishing experiences (Steffen Kernstock); and (in)compatibility of source files (Bernd Raichle). As well, Phil Taylor spoke about the future of TEX and his role as the newly appointed technical director of the NTS project. was used to provide solutions for specialised symbols in a physics context (Torsten Kunz).

Several papers looked at the problems of taking TEX and integrating it with the non-TEX world: a proposal for a standard graphics interface (Friedhelm Sowa); bringing diagrams and graphs from Mathematica and Maple into L^{A}TEX files (Werner Buckhards); and using contents of TEX’s registers in PostScript (Jürgen Glöckner).

Other items on the agenda included a L^{A}TEX3 Question-and-Answer session with Rainer Schöpf, and a description of REdit, a text editor (Michael Rogalla).

Social Events included a guided tour of City Hall on Wednesday, with a reception held by the Mayor of Chemnitz. And on Thursday, we went by bus to Schloß Augustusburg, which houses a famous motorcycle museum; we concluded the evening with dinner at the restaurant there.

Joachim Lammarsch and Rainer Schöpf closed the meeting with the traditional TEX resource talk. The ever-important question to answer: “Where can I get TEX-related software?” File server locations and Internet and Bitnet discussion lists were mentioned, as well as how to use the DANTE TEX Archive at the University of Stuttgart using the mail server or ftp. Last year’s request count and the volume of software retrieved by both mail and ftp were interpreted as eloquent proof as to the usefulness of the Stuttgart Archive.

All in all an interesting and fun meeting I’m glad I attended!

TUG’ 93 at Aston University, UK _{Mimi Burbank mimi@scri.fsu.edu}

There is an air of excitement and anticipation when one goes to the TUG annual meeting; this year was no exception. 1993 was the first year the Annual Meeting was held outside the North American continent, affording many European TEX users to attend — as well as providing economic problems for American users. The program covered topics on system installation, fonts, document design, typography, macros, and several excellent talks on The Future(s) of TEX. There were excellent talks given on languages — Khmer and Catalan — and exciting developments for typesetting music and maps, as well as informative talks on macros and the behaviour of TEX and L^{A}TEX. This year a Bursary Fund was established, with generous contributions made by individuals and organizations, resulting in the attendance by some eastern European users who otherwise could not have attended.

The meeting at Aston was attended by users from all around the globe — Belgium, Canada, Czech Republic, Denmark, Eire, Estonia, France, Germany, Greece, Italy, Israel, Japan, Norway, Poland, Russia, Slovenia, Spain, Sweden, Switzerland, The Netherlands, Ukraine, Ulster, United Kingdom, and the United States.

One highlight of the conference had to be the trip to Stratford to see King Lear at the Royal Shakespeare Theatre — a powerful and most memorable performance. The Shakespeare Library opened its vaults and shared published works dating back to the early 1700s. Speaking as a colonial I must say that the sights, smells, tastes, and sounds of my trip to the UK will remain with me a long time.

The sense of history is quite acute, and reminds one that TUG also has a history, one which continues to evolve and we as members must protect the heritage given us and provide the framework for future users. Networks have brought many of us together electronically; at Aston, many of us met for the first time; many of us argued (and some not for the first time), but the agreement was unanimous that as TEX users, we must continue to evolve.

TUG Board Activities

Summary of Annual Business Meeting TUG ’93 at Aston University, UK

Christina Thiele

President, TEX Users Group

This year’s Annual Business Meeting was held on Thursday, July 29th, from 2:00 to 3:00, chaired by TUG’s president, Christina Thiele. The first part of the meeting was a presentation of various items by members of the Executive Committee, followed by a general question-and-answer session with members attending the Annual Meeting.

Reports The President thanked the TUG ’93 Committee, Peter Abbott for Aston University, the TUG Office staff, and all the attendees for contributing towards a successful and informative meeting.

Peter Abbott, chair of the TUG ’93 Committee, located at Aston University, gave a quick overview of the conference: 9 of 12 courses had been run, and there was a total attendance for both courses and conference of about 168 individuals.

Phil Taylor presented a summary of the results of the Bursary Fund, donations for which had come from many TUG members. A total of some £5,488 had been received, allowing some 18 people from various countries, predominantly Eastern Europe, to attend the first TUG meeting for most of them. On their behalf, Ela Kuczynska (a member of GUST, the Polish user group), read a short statement of thanks.^{1} About £600 remained, and this will go towards the 1994 bursary fund. It should be added here that the money which TUG usually allocates to the Knuth Scholar was redirected to the Bursary Fund, since no scholarship was awarded for 1993.^{2}

1“I would like to express our gratitude to all the persons who gave their donations to the

On behalf of the Executive Director, Pat Monohon, the President provided an update on the TUG office. Of immediate interest and benefit to members was the news that the office staff (Pat, Lisa Ward, and John Berlin) have arranged their hours such that calls can be taken from 3 a.m. to 6 p.m. Pacific Time (!!), and that payments can now be accepted in almost any regularly traded currency — which means it is no longer necessary to always convert to US dollars when sending in membership dues, or payments for other TUG products. Many had already met Pat, Lisa Ward, and John Berlin during the meeting, and members were invited to stop by the office if they were in the Santa Barbara area.

The Treasurer, Bill Woolf, presented an outline of the budget for 1993, which indicated a potential $30,000 deficit. However, since the revenues from the TUG ’93 conference were as yet unknown, nothing more definite could be provided.

Two announcements were made by Barbara Beeton: the first concerned the new release of TEX v.3.1415 (the notice was included in the Thursday edition of the daily TUGly Telegraph newsletter and will also appear in TUGboat). The second was that the deadline for the TTN “A-in-L^{A}TEX” contest has been extended to December 31, 1993.

The President outlined the highlights of the board meeting. The most significant item for members was the issue of the annual meetings. It has been decided that meetings will rotate through three geographic zones: the west coast, then the east coast of North America, and Europe. To start the cycle off, Santa Barbara will be the west coast site for the 1994 TUG Annual Meeting, with a tentative decision for Florida to be the east coast site for 1995. Also with members, the 1994 dues will remain unchanged at the 1993 rates. For the immediate future, TUG’s attention will focus on a membership drive, and on improving and extending services to its members. One further issue of direct impact on TUG members was the upcoming election for board members, which is another step in the transition from an appointed board to an elected one. All 15 positions are open for election; in order to end up with only 5 positions per year coming up for election, the 15 winning candidates will then be selected randomly to hold their seats for one, two, or three years. As well, all terms of office, for both board members and for the president, will begin at the annual meeting: this means future elections will be held in the spring, before the annual meeting.

The President also spoke about the conference at Aston: the excitement of having the meeting in Europe, the opportunity to meet so many new people and to match faces with e-mail addresses, the trip set up to attend the performance of King Lear at Stratford, as well as the special viewing of books arranged at the Royal Shakespeare Library. As always, the success of the meeting lay with the members, both long-time and first-time attendees.

Bursary Fund, as well as to the Bursary Fund Committee, for their kind support. Thanks to them, we have now the possibility to participate in this conference, that would otherwise be impossible for us. It will be a great advantage to our newly-established groups, which have only recently begun to participate in the world-wide TEX community. Thank you very much.”

2See TUGboat 14, #2, p. 102.

Questions and Answers The second half of the Business Meeting was devoted to questions and comments from the audience.

• Nelson Beebe suggested that TUG maintain a forwarding address in perpetuity from the former Providence location to Santa Barbara; it was pointed out that all publications in circulation before 1993 contained the Providence address as contact point for TUG.

Update on this item: Bill Woolf, TUG’s Treasurer, and an Associate Executive Director of the AMS, has said that the former e-mail address of tug@math.ams.org is permanently forwarded to the new address, which is tug@tug.org (as of this month [September] –Ch).

• The problem of delays in receiving TUGboat was raised by Malcolm Clark. He asked if the scheduled vs. actual dates of publication for the past 5 years might be assembled, to determine if current delays were no different from those in the past, or if they had become worse.

The President’s response was that this was a serious issue, and that it was clear that recent delays were indeed worse. She pointed out that finding alternative production solutions for a publication which had 14 years of history could not be found overnight. She has been discussing other options with the TUGboat editor, Barbara Beeton, since early spring and had made it one of her priority items while in office.

Phil Taylor rose later on to offer effusive praise and thanks for the work done by the editor, and this was greeted with a generous round of applause by the audience, to the editor’s embarrassment.

• The question of BibTEX and when it might be available was raised. The President responded that she’d been in contact with Oren Patashnik, who felt that it could not be finalised before the end of this year. As well, there were complications arising as the L^{A}TEX3 project tried to ensure that there would be harmony between BibTEX and the new L^{A}TEX.

• Jonathan Fine rose to suggest that authors of public domain software consider having TUG serve as distributor of their wares, and perhaps even have TUG involved in locating funds to support writers of PD software. The President suggested that Jonathan serve as the collecting point of information and comment on this issue, and present a report to TUG.

• An additional item raised by Malcolm was that of having an Index to TUGboat. Barbara replied that an Index had been in progress for quite some time; progress was slow, but she could report that the issues up to volume 8 were now done. As with many volunteer efforts, time was the major constraint. However, all TUGboat tables of contents have been installed on the CTAN archives, along with the style file needed to typeset them. As well, a KWIC index generated from the titles in them is also present in the archives, thanks to Nelson Beebe.^{3}

• The suggestion was made, based on a presentation by the editors of the Polish newsletter GUST, that the TUGboat macros might also benefit from being recast in WEB. The editor was not sure this was feasible, but said she would investigate the issue further.

• And finally, it was suggested that one way of helping the editorial and production work on TUGboat move more smoothly and rapidly, would be for authors submitting items for publication to take a little extra time to ensure their submissions were complete (check all \input statements to ensure all additional macro files were included), that their file runs on a different machine, especially with respect to non-standard or non-CM fonts; check spelling; etc. Having a colleague read over the submission, as well as test run it, would also help debug some of the more obvious errors, all of which can add to the editor’s task of getting material ready for publication on time.

The meeting ended at 3 p.m., in time for the afternoon break.

Notices

• 1993 Board Election Results _{Joachim Lammarsch Chair, Election Committee}

The terms of all fifteen members of the first elected Board of Directors will expire on December 31, 1993. The election to choose the new Board members was scheduled for this fall and nominations of all members were invited.

For this election only, the length of some terms has been altered so that not all members of the Board are replaced at the same time, but only onethird. For Board members elected this year, all terms will begin January 1, 1994. Five members of the new Board will have terms that end with the 1995 annual meeting, five with the 1996 annual meeting and five with the 1997 annual meeting. The members of each of the three groups have been chosen by lot by the TUG acountant.

The deadline for nominations was September 1, 1993 and the TUG office received twelve nomination forms. Unfortuately one was not complete and had to be declared invalid.

3Details appear in TUGboat 14, #2, p. 101.

Therefore it was not necessary to mail ballots to all members. In accordance with the Election Procedures all eleven candidates were declared as elected. In accordance with the bylaws our president will appoint four other persons to fill the four remaining vacancies.

The new members of the Board of Directors are (selected by term):

3.5-year term  2.5-year term  1.5-year term
Luzia Dietsche  Mimi Burbank  Barbara Beeton
George Greenwade  Jackie Damrau  plus four people,
Yannis Haralambous Michael Ferguson to be named.
Jon Radel  Peter Flynn
Sebastian Rahtz  Nico Poppelier

As chair of the Election Committee I would like to express my thanks to the candidates and all the other people who helped to organize the election.

However, if we are not able to find more than fifteen candidates I feel it’s time to consider: Is an election by such a democratic method useful? Would it be better to return to appointing members to the Board of Directors? Are there other solutions? I would be glad to receive suggestions from our members.

• Technical Council:

Special Interest Working Groups (SI-TWGs)

Michael Ferguson Chair, Technical Council

The primary purpose of the TEX User Group’s Technical Council is to provide a forum and mechanism for studying technical issues concerning TEX, , and their auxilliary support systems. Although this includes such important service, standardisation, future tools projects,^{1} it is the “Special Interest Working Groups (SI-TWG)” that will have the most interest for the ordinary TEX user. SI-TWGs are dedicated to the exploitation of TEX systems for the specialised needs of particular user groups. These vertically integrated interest groups are destined to be an incredibly exciting part of the TEX community.

Our first SI-TWG was “TEX for the Disabled” with T.V. Raman as chair (raman@cs.cornell.edu). Raman, who is blind, has created a very impressive system for interpreting the control codes and structures of TEX by means of audio tones. His TWG is not only concerned with making TEX more accessible for the blind, but also for persons with other disabilities. This group has been able to connect people who need and people who produce such things as braille fonts for mathematics, Computerized Books for the Blind, and appropriate online versions of the L^{A}TEX manual. There has also been discussion and offers within the group to write some specialized programs, such as adapting DOS for the vision impaired. This SI-TWG is an exciting beginning to, what I believe, will be a very important part of the TEX environment.

1Examples of such projects include Archives (TWG-Archives: George Greenwade, TWG-Chair), Multiple Language Coordination (TWG-MLC: Yannis Haralambous, TWG-Chair), and Independent Research Projects (IRP-TWG: Alan Hoenig, TWG-Chair) which includes the L^{A}TEX3 project (Frank Mittelbach, Project Chair). See TTN 1,3:5–8 for details.

At this moment, we are in the process of organizing two new SI-TWGs. The first is to address the needs of those that are using, need to use, or merely interested in, the coupling of TEX and SGML (contact Ken Dreyhaupt at dreyhaupt@spint.compuserve.com). This is a very important area for the enhancement of cooperation and understanding between two communities, and will also serve as a means of discussion and research for those that must straddle both areas. The other new SI-TWG being organized concerns the relation of TEX and Adobe’s new file format Acrobat (Doug Henderson, drh@bluesky.com).

SI-TWGs may reflect the specialized technical interests of a particular community such as publishers, scientific organizations, chemists, astronomers, physicists, historians, logicians, theoretical computer scientists, or musicians, or perhaps satisfy a need for the possibly specialized use of a tool such as L^{A}TEX. One particular area might be Human/Computer interfaces for L^{A}TEX.

What does it take to start an SI-TWG? First, and foremost, it takes an enthusiastic person with a burning interest. There also should be a user community with similar interests that will come together, usually electronically, for discussions and actions. Each TWG has a chair and a member of the Technical Council, who is sometimes the chair, to act as a liaison between the TWG and the TUG Board. The TWG chair is expected to report, usually through the Technical Council Liaison, its activities to the Technical Council and hence the board. The chair is also encouraged to report on the TWG’s activities in any forum that he/she feels beneficial and to seek cooperation with individuals and institutions as appropriate. If there is any action that a TWG would like to report to the board for formal (TUG) recognition, this action or recommendation will be evaluated by the Technical Council, and then brought to the board. It is anticipated that there will be little need for such action by a SI-TWG but the possibility is there.

The TWG chair is completely responsible for the membership and activities of the TWG. The Technical Council and its liaison is charged only with reporting the activities of the TWG and carrying recommendations to the TUG board. If you would like to participate in a particular TWG, please contact the Chair. If you would like to start a TWG, please contact any member of the Technical Council.

       Michael Ferguson (Chair) mike@inrs-telecom.uquebec.ca
       Yannis Haralambous  yannis@gat.citilille.fr
       Alan Hoenig  ajhjj@cunyvm.cuny.edu

• Conference Planning Committee_{Jackie Damrau Chair, Conference Planning Committee}

Following recommendations from our committee, the TUG board has approved a geographic rotation through three zones: North America — West Coast, North America — East Coast, and Europe. This approach allows TUG to reach many of its members by giving all an opportunity to attend an annual meeting periodically without undue expense.

The TUG board has approved Santa Barbara, California as the 1994 annual meeting site. Plans are in progress by the local Santa Barbara TEX Users Group to offer enticing accommodations, meeting arrangements, and a pleasant place to meet with other TUG members from near and far.

We are now soliciting the membership for volunteers to host future annual meetings for 1995–1997. Keeping with our geographical scheme, we are looking for offers from North America — East Coast users (1995); European users (1996); and North America — West Coast users (1997). The committee has received tentative offers from Florida (St. Petersburg, Florida, has been accepted as a candidate site) and Ireland. It is never too late to offer your campus or institution as a possible meeting site.

What do you need to do to offer your site? Using the recorded experience of past conference organizers, the committee has compiled a short checklist for volunteers to complete and return for an initial bid to host a meeting. Along with this checklist, the committee provides organizers with a summary of tips — sort of a “things to do” and “things to avoid”.

Would you like to host an annual meeting at your site for 1995, 1996, or 1997? If so, please contact committee chair, Jackie Damrau:

Jackie Damrau, Chair  Phone: 214-708-6048
TUG Conference Planning Committee FAX: 214-708-5143

 P.O. Box 875  E-mail: tugcpc-l@irlearn.bitnet

Red Oak, Texas 75154-0875, USA

1994 Annual Meeting

31 July–4 August, 1994

Santa Barbara, California

TEX Users Group 1993/1994 Course Schedule

Beginning/Intermediate TEX* Intensive Course in L^{A}TEX*
Santa Barbara Feb. 7–11  Santa Barbara Jan. 31–Feb. 4
Modifying L^{A}TEX Style Files* Adv’d TEX and Macro Writing*
Santa Barbara Feb. 28–Mar. 4  Santa Barbara Nov. 1–5
Santa Barbara Feb. 14–18

Practical SGML and TEX* Publishers*

Santa Barbara Nov. 8–9

*Lab classes — computers will be provided for all students

• TUG courses are small, with 8–15 students in most classes.

• Dates and locations subject to change. Direct course inquiries to John Berlin at the TUG Office. Anyone interested in courses dealing with TEX and SGML should contact John as well (john@tug.org).

• TUG office: phone 805-963-1338 FAX 805-963-8358 tug@tug.org

On-Site Courses in TEX and L^{A}TEX from TEX Users Group

• Courses in TEX, L^{A}TEX, SGML and TEX, PostScript, or TEX for Publishers tailored to the needs of your group

• Courses at every level from beginning to advanced

• Five full days of instruction at your site

• One-week course fee includes all instructor fees and expenses plus textbooks and other materials for up to 15 students

Upcoming Events

3 Nov.

GUTenberg: “Journée pour l’utilisateur TEX” and General Meeting. ENS, salle Dussane, 45 rue d’Ulm, Paris 5ème.

                            journee-gut@ens.fr

18 Nov.

NTG 12th Meeting: “(L^{A})TEX User Environment”. Den Bosch, OCÉ.

Gerard van Nes

                            vannes@ecn.nl

18–19 Nov.

“Jump Starting Your SGML Implementation” (seminar). Goudy Centre, Rochester Inst. of Technology, New York.

Charlene O’Hanlon Tel.: +1 716-475-4951

29 Nov.

Nordic Group: Helsinki University, Finland.

Martti Nikunen

                            nikunen@cc.helsinki.fi

16–18 Feb.

DANTE’94:

10th general meeting, Münster, Germany.

Wolfgang Kaspar

                            kaspar@dmswwu1a.
                            uni-muenster.de

11–15 April Joint Conference: RIDT’94,

EP94, TEP’94, PODP94.

Darmstadt, Germany.

                            ep94@gmd.de,
                            ridt94@irisa.fr

Spring 1994 NTG 13th Meeting: “(L^{A})TEX, , and tools education.” Groningen, at RUG.

Gerard van Nes

                            vannes@ecn.nl

{31 July– \atop 4 Aug.}

TUG ’94:

15th Annual Meeting, Santa Barbara, California.

Debbie Ceder

                            tug94@tug.org

Note: Also consult the “Calendar” in the previous issue of TUGboat for more dates and details.

In Memoriam

Yuri Melnichuk

Yuri Melnichuk, a reader in computing mathematics at Lviv Polytechnical Institute, in the Ukraine, and one of the participants at the Aston TUG ’93 conference and courses, died suddenly of a heart attack while at the University of York on Friday, August 13th.

Yuri was a fairly frequent visitor to York, where he was working on a joint book on number theory with Maurice Dodson. During his time at York he had been introduced to TEX, and with the help of other colleagues in the Mathematics department there had developed this interest. Realising its potential value to the academic community in the Ukraine he had contacted others, with a view to establishing good links between institutions and individuals.

His concerns were not just mathematical: he was also active in ensuring that the British Council and the IEE were in contact with relevant bodies in the Ukraine. Just before he died he was searching the net for a PC version of Ada to take back for computer scientists in his institution.

Last year, he had been instrumental in beginning a TEX users group in Lviv and was beginning to coordinate its activities over the whole country.

He was a dynamic personality, with infectious vision and enthusiastic plans. An excellent and hospitable host, his good humour was matched with determination. His ability to bring people together was a tribute to his vision of cooperation and his own engaging nature. His loss is a blow to the many friends he had made, to his colleagues in the Ukraine, and to the TEX community worldwide.

Malcolm Clark

TTN Schedule for Submissions

Jan. issue (1994) Dec. 1st
Apr. issue  March 15th
July issue  June 1st
Oct. issue  Sept. 15th

1993 TTN Index

Notation: All references show the volume,number:page; outside back cover is noted as C4 (Index prepared by Jackie Damrau).

– A –

The A-in-L^{A}TEX Contest 2,2:28; 2,4:14 Archives

UK TEX archive on CD-ROM 2,3:14 The CTAN archive 2,3:15

An Author’s Ps@lm 2,4:28

– C –

1993 Course Schedule 2,1:24; 2,2:26; 2,3:27;

2,4:24 Conference Planning Committee 2,3:26; 2,4:22

– D –

Did you know … ? 2,1:2; 2,2:2

– E –

Editorials 2,1:1; 2,2:1; 2,3:1; 2,4:1 Eight Great Reasons to Join the TEX Users Group 2,3:C4 Board Elections 2,2:25; 2,3:23; 2,4:20

Nomination form 2,3:24

– F –

Feature Articles International quotations 2,3:5

– H –

“Hey — it works!”

Correct spacing for tables and arrays 2,3:10 Experiences with tables 2,1:7 Footnotes with verbatim material 2,4:9 Headings in reference lists 2,2:8;

Old vs. new L^{A}TEX problem 2,3:9 \howto pattern match in TEX 2,2:10 Labelling equations by name 2,3:12 Lining up decimal points 2,2:7 More on margin marks 2,3:9 1-column material below 2-column 2,4:6 Striking through text 2,4:7 Thicker \hlines in L^{A}TEX alignments 2,4:8

– I –

In Memoriam

Yuri Melnichuk 2,4:26 Index 2,4:27 International quotations 2,3:5

– L –

(L^{A})TEX News Announcing OzTEX 1.5 2,3:13 BibTEX Hypercard stack 2,3:13 Bits and pieces … 2,4:11 DC fonts — Q and A (II) 2,1:10 Files to print up font tables 2,1:13 Jones Index of TEX Macros 2,2:12 The CTAN archives 2,3:15 TTN electronic distribution 2,1:13 TTN index 2,2:13 UK TEX archive on CD-ROM 2,3:14

L^{A}TEX3 Project Volunteer Work 2,2:3 L^{A}TEX3 Update: Coming Soon! 2,4:11

– N –

New Publications 2,1:6; 2,2:6; 2,3:4; 2,4:3 News from Around …

La Lettre GUTenberg 2,2:20 News from GUTenberg 2,1:15 NTG happenings 2,2:19 TEX waves from Estonia 2,2:17 uk → ue 2,1:14

Nomination form 2,3:24 Notices

Technical Council 2,3:25; 2,4:21 Conf. Planning Committee 2,3:26; 2,4:22 Election Committee 2,4:20

– O –

Other TEX User Groups 2,1:27

– P –

Positions Available 2,4:2

– R –

Reports on Meetings 6th Annual Nordic meeting (Nov. ’92) 2,2:21 BoF session at TUG’92 2,1:16 DANTE’93 in Chemnitz (March ’93) 2,4:15 Joint Conferences (April ’94) 2,1:22 News about CyrTUG … 2,1:17 NTG meeting (Nov. ’92) 2,1:16 NTG’s Lustrum meeting (June ’93) 2,3:22 TUG’93 2,1:20; 2,1:C4; 2,2:C4; 2,3:19,

2,4:16

News Updates 2,2:23; 2,3:19 UKTUG (April ’93) 2,1:19

– T –

TUG Board Activities Board Elections 2,2:25; 2,3:23; 2,4:20 Summary of Annual Business Meeting 2,4:17 Message from New President 2,1:23 Technical Council 2,3:25; 2,4:21 TUG Committees

Conf. Planning Comm. 2,3:26; 2,4:22 Election Committee 2,4:20

Typographer’s Inn 2,1:3; 2,2:4; 2,3:3; 2,4:4

– U –

Upcoming Events 2,1:25; 2,2:27; 2,3:28; 2,4:25 User Groups (list of) 2,1:27

– Y –

“You Rang … ?”

All about ACT 2,4:12

How can I get an indent after a sectioning

command? 2,3:16 How do I change the line spacing? 2,3:18 How do I get an abstract in the right place in L^{A}TEX using \maketitle? 2,3:17 Q&A: How do I print envelopes? 2,2:14 What is this column? 2,2:14

An Author’s Ps@lm

There are times when an author may find it difficult to justify the choice of a particular word, style of syntax, or punctuation, yet know in his or her h[e]art that a justification could be found and adduced. In those circumstances, the author may wish to seek the assistance of a Greater Power: I find the following very reassuring:

Hart is my shepherd, I shall not want; He maketh me to punctuate correctly; He restoreth my missing apostrophes. Yea, though I walk through the valley of the shadow of Chicago I shall fear no evil; For He is with me; He annointeth my head with sound practices; My \hboxes runneth over. Surely \goodness shall follow me all the days of my life, And I will dwell in the house of TEX forever.

Philip Taylor

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Positions Available . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

New Publications Peter Schmitt . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Typographer’s Inn Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

“Hey — it works!”

Jeremy Gibbons . . . . . . . . . . . . . . . . . . . . . . . . . . One-column material below two-column Ch. Thiele . . 6 Striking through text David Lau . . . . . . . . . . . . . 7 Thicker \hlines in L^{A}TEX alignments Anita Z. Hoover . 8 Footnotes with verbatim material Jeremy Gibbons … 9

(L^{A})TEX News

L^{A}TEX2ε— A New Version of L^{A}TEX

Chris Rowley and Frank Mittelbach . . . . . . . . . . . . 10 Bits and pieces …

Christina Thiele . . . . . . . . . . . . . . . . . . . . . . . 11

You Rang …?

Ed Baker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

A-in-L^{A}TEX Contest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

Reports on Meetings DANTE’93 in Chemnitz . . . . . . . . . . . . . . . . . . . . . 15 TUG ’93 at Aston University . . . . . . . . . . . . . . . . . . . 16

TUG Board Activities Summary of Business Meeting (Aston) . . . . . . . . . . . . . 17 Notices

1993 Board Election Results . . . . . . . . . . . . . . . . . 20 Technical Council . . . . . . . . . . . . . . . . . . . . . . . 21 Conference Planning Committee . . . . . . . . . . . . . . 23

TEX Users Group 1993 Course Schedule . . . . . . . . . . . . . . . . . 24

Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

In Memoriam: Yuri Melnichuk . . . . . . . . . . . . . . . . . . . . . . 26

1993 TTN Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Volume 2, No. 4, 1993

TUG’94 Santa Barbara, California 31 July – 4 August

Call for Papers

The 15th Anniversary Meeting of the TEX Users Group will be held in Santa Barbara, California from 31 July through 4 August 1994. Those wishing to present papers must have their titles and outlines submitted to the program committee at tug94@tug.org by February 1, 1994, completed papers by May 20, 1994.

Since TEX and applications and interests are as varied as our users, we are encouraging papers over the entire range of related topics, with a particular focus on innovation. Let us take a fresh look at what we have and envision new areas of use.

Open Position American Mathematical Society

Technical Support Department

The AMS is seeking an individual to join six other TEX programmers in the Society’s Technical Support Department. The successful candidate will possess a bachelor’s degree in C.S. (or equivalent), a minimum of three years’ experience in the publishing industry, and all or most of the following qualifications:

• Advanced TEX macro writing

• C programming language

• Experience in electronic prepress techniques (electronic halftones, imposition, etc.)

• VMS, Unix, Mac-OS, DOS, and Windows operating systems • Knowledge of the PostScript

language

• Experience with traditional desktop publishing on Macintosh computers

• Acute attention to detail, strong verbal and organizational skills

The Society is an equal opportunity employer and provides excellent benefits. The building is wheel-chair accessible. Interested applicants should send résumé and salary requirements to: Director of Human Resources, American Mathematical Society, P.O. Box 6248, Providence, Rhode Island 02940.